# Broadway: Boost Your Text-to-Video Generation Model in a Training-free way

### v1

@@ Abstract
The text-to-video (T2V) generation models, offering convenient visual creation, have recently garnered increasing attention. Despite their substantial potential, the generated videos may present artifacts, including structural implausibility, temporal inconsistency, and a lack of motion, often resulting in near-static video. In this work, we have identified a correlation between the disparity of temporal attention maps across different blocks and the occurrence of temporal inconsistencies. Additionally, we have observed that the energy contained within the temporal attention maps is directly related to the magnitude of motion amplitude in the generated videos. Based on these observations, we present BroadWay, a training-free method to improve the quality of text-to-video generation without introducing additional parameters, augmenting memory or sampling time. Specifically, BroadWay is composed of two principal components: 1) Temporal Self-Guidance improves the structural plausibility and temporal consistency of generated videos by reducing the disparity between the temporal attention maps across various decoder blocks. 2) Fourier-based Motion Enhancement enhances the magnitude and richness of motion by amplifying the energy of the map. Extensive experiments demonstrate that BroadWay significantly improves the quality of text-to-video generation with negligible additional cost.
Figure 1: Given a diffusion based text-to-video (T2V) backbone, BroadWay can improve its synthesis quality in a training-free and plug-and-play manner, enhancing both the temporal consistency and the motion magnitude in the sampled results.
@@ 1 Introduction
In recent years, the field has observed substantial progress in the evolution of diffusion-based models specifically dedicated to video generation tasks, notably in text-to-video synthesis (Khachatryan et al., 2023; Blattmann et al., 2023b; Guo et al., 2023b; Chen et al., 2024). Despite these advancements, the practical applicability of generated videos remains limited due to inadequate quality. This suboptimal performance is characterized by two predominant issues: firstly, a portion of the generated videos exhibit structurally implausible and temporally inconsistent artifacts, and secondly, another subset of the generated videos demonstrates markedly restricted motion, bordering on the static nature of a still image. Prior methodologies have primarily concentrated on enhancing video generation quality through advances in training mechanisms, such as improving the quality of training data (Blattmann et al., 2023a), scaling training data (Wang et al., 2024b), refining model architecture (Hong et al., 2022) and training strategies (Chen et al., 2024). However, these approaches often entail substantial costs. This work endeavors to improve video generation quality in the inference phase, specifically in the realm of text-to-video generation, without necessitating training, introducing additional parameters, augmenting memory or sampling time.
Figure 2: Generated videos with richer motion typically exhibit a higher energy.
@@
In current video generation models, an encoder-decoder architecture (Ronneberger et al., 2015) is typically utilized, wherein the decoder is comprised of multiple blocks. Each block integrates several temporal attention modules (Guo et al., 2023b), facilitating the modeling of motion within the generated videos. We have two observations about the temporal attention module. The first is a correlation between artifact presence and the inter-block divergence of temporal attention maps. Specifically, video generation processes exhibiting structurally implausible and temporally inconsistent artifacts demonstrate greater disparity between the temporal attention maps of different decoder blocks. Conversely, processes devoid of such evident artifacts exhibit reduced disparity among these maps, as illustrated in Fig. 3(a). The second is a correlation between the amplitude of motion in generated videos and the energy of the corresponding temporal attention maps, defined in the method section. Specifically, videos that exhibit a higher degree of motion amplitude and a richer variety of motion patterns are observed to possess greater energy within their temporal attention maps, as illustrated in Fig. 2.
@@
Based on the observations, we present BroadWay, a training-free approach with negligible additional cost to improve the generation quality of T2V diffusion models. BroadWay is composed of two principal components: Temporal Self-Guidance and Fourier-based Motion Enhancement, both meticulously engineered to refine the temporal attention module within T2V models. Temporal Self-Guidance leverages the temporal attention map from the preceding block to inform and regulate that of the current block. This approach effectively mitigates the disparity between the temporal attention maps across various decoder blocks, thereby normalizing their disparity. As a result, videos that initially exhibit structural implausibility and temporal inconsistency, significantly reduce such artifacts through the application of Temporal Self-Guidance, as shown in the first and second rows in Fig. 1. Furthermore, Fourier-based Motion Enhancement modulates the high-frequency components of the temporal attention map, thereby amplifying the energy of the map, as detailed in the methodology section. This enhancement circumvents the generation of videos that closely resemble static image. With the Fourier-based Motion Enhancement, videos that were previously characterized by minimal motion exhibit an increased amplitude and a more diverse range of motion patterns, as illustrated in the third and last rows in Fig. 1.
@@
We evaluate the performance of BroadWay on various popular T2V backbones, including those with additional motion modules trained from frozen T2I models and those trained end-to-end directly for T2V tasks. Our experiments show promising results, demonstrating the effectiveness and strong adaptability of BroadWay. Furthermore, additional experiments reveal that BroadWay also exhibits potential in the image-to-video (I2V) domain, further expanding the applicability of BroadWay across various video generation tasks.
@@
Our contributions are summarized: (1) We conduct a deeper analysis of the temporal attention module widely adopted in current T2V backbones, and observe two correlations between the generated videos and corresponding temporal attention maps. (2) We propose BroadWay, which significantly improves the quality of text-to-video generation without necessitating training, introducing additional parameters, augmenting memory or sampling time. (3) BroadWay can be seamlessly integrated with various mainstream open-source T2V backbones like AnimateDiff and VideoCrafter2, demonstrating strong applicability and extensibility.
Figure 3: MSE distance between temporal attention maps of different levels in UNet.
@@ 2.1 Text-to-Video Diffusion Models
Given a textual prompt, text-to-video (T2V) diffusion models (Singer et al., 2022; Hong et al., 2022; Wang et al., 2023b; Chen et al., 2023a; Wang et al., 2023a; 2024a; Khachatryan et al., 2023) aim to synthesize image sequences that maintain both temporal consistency and textual alignment. Unlike text-to-image (Ding et al., 2021; Zeqiang et al., 2023; Saharia et al., 2022; Podell et al., 2023) that emphasizes perfecting individual images, T2V poses a heightened challenge of maintaining both visual aesthetics for each frame and the realistic motion between frames. To this end, most approaches incorporate extra motion modeling modules into existing image diffusion architecture, leveraging the underlying image priors. For instance, AnimateDiff (Guo et al., 2023b) introduced trainable temporal attention layers to frozen text-to-image models to effectively capture the frame-to-frame correlations Some works (Blattmann et al., 2023b; Chen et al., 2024) combined temporal convolution modules and temporal attention layers for modeling short/long range dependencies. To alleviate motion synthesis difficulty, Ge et al. (Ge et al., 2023) suggested employing temporally related noise to enhance temporally consistent. Nevertheless, due to the scarcity of high-quality video data and the intricacies of motion synthesis, the current available T2V models still struggle to harmonize motion strength with motion consistency. This work identifies that the consistency across temporal attention blocks indicates the continuity of synthesized video sequences while the energy within the temporal attention maps dominates the magnitude of motion, and thus proposes a training-free strategy to unlock the potential of exiting T2V models by encouraging uniform motion modeling and enhanced frequency energy.
@@ 2.2 Diffusion feature control
Controlling targeted diffusion features to manipulate specific attributes has been demonstrated to be an effective strategy in the realm of image and video synthesis (Chefer et al., 2023; Kim et al., 2023b; Xiao et al., 2023; Liu et al., 2023; Qi et al., 2023),. Prompt2Prompt (Hertz et al., 2022) revealed that the cross attention maps domain the image layout. DSG (Yang et al., 2024) proposed that spatial means of diffusion features represent the appearance, which offers simple approach for image property manipulation, such as size, shape, and location. FreeControl (Mo et al., 2023) suggested to perform image structure guidance by aligning the PCA features with given reference image in spatial self-attention block, providing a versatile counterpart of ControlNet (Zhang et al., 2023). DIFT (Tang et al., 2023) observed that the semantic corresponding can be directly extracted by spatially measuring the difference between diffusion feature. MotionClone (Ling et al., 2024) demonstrates the sparse control of temporal attention maps facilitates a training-free motion transfer, enabling reference-based video generation. FreeU (Si et al., 2024) suggested re-weighting the contribution of skip features and backbone features by using spectral modulation and structure-related scaling, promoting the emphasis on backbone semantics. In this work, we propose Temporal Self-Guidance to facilitates uniform motion modeling across blocks by narrowing the disparities between temporal attention maps. This is work together with Fourier-based Motion Enhancement, which boosts motion magnitude by amplifying frequency energy, thus elevating the quality of the generated videos
@@ 3 Preliminary
In the context of T2V generation, latent diffusion model (Rombach et al., 2022b) is widely as backbone as its significant advancement in image synthesizing. Typically, based on a pre-trained autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$, video sequences are projected into the latent space, in which a denoising network $\epsilon_{\theta}$ is encouraged to learn the mapping from noised video latent $z_{t}$ to pure video latent $z_{0}$. Mathematically, the noised video latent $z_{t}$ obeys the following distribution:
@@
where $\bar{\alpha}_{t}$ is a pre-defined parameter representing noise schedule (Ho et al., 2020), $\epsilon\sim\mathcal{N}(0,1)$ is the added noise, and $t\sim\mathcal{U}(1,T)$ denotes time step. To restore $z_{0}$ from $z_{t}$, denoising network $\epsilon_{\theta}$ is forced to estimate the noise component in $z_{t}$, which can be expressed as:
@@
where $c$ represents the textual prompt of $z_{0}$. During sampling, $z_{t}$ is initialized with Gaussian noise and undergoes iterative denoising conditioned on $c$ for prompt-aligned generation.
@@ 4.1 Temporal Self-Guidance
Temporal attention modules are extensively integrated at various hierarchical stages within the upsampling blocks of T2V architectures (Blattmann et al., 2023b; Guo et al., 2023b; Chen et al., 2023a; 2024). These modules, derived from different tiers of the diffusion UNet, are employed to capture inter-frame dependencies at multiple resolutions. Although the multi-level progressive refinement approach in modeling frame-wise correlations offers advantages, our observations indicate that the temporal attention maps across different hierarchical levels can exhibit considerable discrepancies, potentially leading to structurally implausible or temporally inconsistent video outputs. To substantiate this hypothesis, we analyzed 100 structurally and motion-degraded videos alongside 100 well-generated videos. We computed the mean and standard deviation of the mean squared error (MSE) distances between the temporal attention maps of up_blocks.1 and subsequent blocks as illustrated in Fig. 3 (a). Our findings reveal that significant disparities between temporal attention maps across different blocks are associated with the occurrence of implausible structures and temporal inconsistencies in the generated videos.
To mitigate the excessive divergence between temporal attention maps across various upsampling blocks, we introduce a straightforward yet potent temporal self-guidance mechanism. This mechanism involves the infusion of the temporal attention map of up_blocks.1 into subsequent blocks, modulated by a guidance ratio $\alpha$. The adjustment is mathematically represented as:
@@
where $\mathcal{A}_{m}$ denotes the temporal attention map of the $m$-th upsampling block ($m=2,3$), and $\mathcal{A}_{1}^{m}$ refers to the temporal attention map of up_blocks.1, which is upsampled to match the spatial dimensions of $\mathcal{A}_{m}$. As depicted in Fig. 3 (b) and Fig. 4, the implementation of temporal self-guidance effectively alleviates the excessive modeling disparity between different hierarchical levels of temporal attention modules, thereby diminishing structurally implausible and temporally inconsistent artifacts in the resultant video generation.
@@
Figure 4: Temporal Self-Guidance. Temporal Self-Guidance contributes to the restoration of collapsed structures and consistency of motion in the synthesized video.
@@ 4.2 Fourier-based Motion Enhancement
$E=\frac{1}{F}\sum_{i=0}^{F-1}\sum_{j=0}^{F-1}||\mathcal{A}_{...,i,j}||^{2},$ (4)
as illustrated in Fig. 5 (a). To substantiate the correlation between the energy of the temporal attention map and the motion magnitude in the generated video, we employ the RAFT (Teed & Deng, 2020) to extract the optical flow, using the average magnitude of this flow as a metric for motion strength. Our findings reveal a positive correlation: videos with greater motion magnitudes are associated with higher energies within their temporal attention maps. This insight motivates us to manipulate the motion magnitude in the generated videos by modulating the energy intensity of the temporal attention maps. By doing so, we aim to enhance the dynamism and variability of the motion in the videos.
Figure 5: Visualization of energy in temporal attention map. (a) The energy map of the generated video. (b) Videos with larger motion magnitude typically exhibit higher energy, where the motion magnitude is calculated using the optical flow of the generated videos.
4.2.2 motion enhancement by frequency spectrum re-weighting

### v3

# ByTheWay: Boost Your Text-to-Video Generation Model to Higher Quality in a Training-free Way

@@ Abstract
The text-to-video (T2V) generation models, offering convenient visual creation, have recently garnered increasing attention. Despite their substantial potential, the generated videos may present artifacts, including structural implausibility, temporal inconsistency, and a lack of motion, often resulting in near-static video. In this work, we have identified a correlation between the disparity of temporal attention maps across different blocks and the occurrence of temporal inconsistencies. Additionally, we have observed that the energy contained within the temporal attention maps is directly related to the magnitude of motion amplitude in the generated videos. Based on these observations, we present ByTheWay, a training-free method to improve the quality of text-to-video generation without introducing additional parameters, augmenting memory or sampling time. Specifically, ByTheWay is composed of two principal components: 1) Temporal Self-Guidance improves the structural plausibility and temporal consistency of generated videos by reducing the disparity between the temporal attention maps across various decoder blocks. 2) Fourier-based Motion Enhancement enhances the magnitude and richness of motion by amplifying the energy of the map. Extensive experiments demonstrate that ByTheWay significantly improves the quality of text-to-video generation with negligible additional cost.
Figure 1: Unlock the potential of pretrained text-to-video (T2V) generation models in a training-free approach. (1) ByTheWay helps to enhance structural plausibility and temporal consistency in generated videos, significantly reducing artifacts and flickering. (2) ByTheWay contributes to enriching motion patterns and amplifying the motion magnitude in generated videos. Further, ByTheWay can be seamlessly integrated into various powerful T2V backbones (e.g., AnimateDiff[9] and VideoCrafter2[7]) in a plug-and-play manner, serving as a highly extensible module without introducing additional parameters or sampling cost.
00footnotetext: * indicates equal contribution, † indicates corresponding author
@@ 1 Introduction
In recent years, the field has observed substantial progress in the evolution of diffusion-based models specifically dedicated to video generation tasks, notably in text-to-video synthesis [58, 8, 9, 7]. Despite these advancements, the practical applicability of generated videos remains limited due to inadequate quality. This suboptimal performance is characterized by two predominant issues: firstly, a portion of the generated videos exhibit structurally implausible and temporally inconsistent artifacts, and secondly, another subset of the generated videos demonstrates markedly restricted motion, bordering on the static nature of a still image. Prior methodologies have primarily concentrated on enhancing video generation quality through advances in training mechanisms, such as improving the quality of training data [36], scaling training data [69], refining model architecture [64] and training strategies [7]. However, these approaches often entail substantial costs. This work endeavors to improve video generation quality in the inference phase, specifically in the realm of text-to-video generation, without necessitating training, introducing additional parameters, augmenting memory or sampling time.
@@
In current video generation models, an encoder-decoder architecture [57] is typically utilized, wherein the decoder is comprised of multiple blocks. Each block integrates several temporal attention modules [9], facilitating the modeling of motion within the generated videos. We have two observations about the temporal attention module. The first is a correlation between artifact presence and the inter-block divergence of temporal attention maps. Specifically, video generation processes exhibiting structurally implausible and temporally inconsistent artifacts demonstrate greater disparity between the temporal attention maps of different decoder blocks. Conversely, processes devoid of such evident artifacts exhibit reduced disparity among these maps, as illustrated in Fig. 2(a). The second is a correlation between the amplitude of motion in generated videos and the energy of the corresponding temporal attention maps, defined in the method section. Specifically, videos that exhibit a higher degree of motion amplitude and a richer variety of motion patterns are observed to possess greater energy within their temporal attention maps, as illustrated in Fig. 2(c).
@@
Based on the observations, we present ByTheWay, a training-free approach with negligible additional cost to improve the generation quality of T2V diffusion models. ByTheWay is composed of two principal components: Temporal Self-Guidance and Fourier-based Motion Enhancement, both meticulously engineered to refine the temporal attention module within T2V models. Temporal Self-Guidance leverages the temporal attention map from the preceding block to inform and regulate that of the current block. This approach effectively mitigates the disparity between the temporal attention maps across various decoder blocks, thereby normalizing their disparity. As a result, videos that initially exhibit structural implausibility and temporal inconsistency, significantly reduce such artifacts through the application of Temporal Self-Guidance, as shown in the first and second rows in Fig. 1. Furthermore, Fourier-based Motion Enhancement modulates the high-frequency components of the temporal attention map, thereby amplifying the energy of the map, as detailed in the methodology section. This enhancement circumvents the generation of videos that closely resemble static image. With the Fourier-based Motion Enhancement, videos that were previously characterized by minimal motion exhibit an increased amplitude and a more diverse range of motion patterns, as illustrated in the third and last rows in Fig. 1.
@@
We evaluate the performance of ByTheWay on various popular T2V backbones, including those with additional motion modules trained from frozen T2I models and those trained end-to-end directly for T2V. Our experiments show promising results, demonstrating the effectiveness and strong adaptability of ByTheWay. Moreover, experiments reveal that ByTheWay also exhibits potential in the image-to-video (I2V) domain, further expanding the applicability of ByTheWay across various video generation tasks.
@@
Our contributions are summarized: (1) We conduct a deeper analysis of the temporal attention module widely adopted in T2V backbones, and observe two correlations between the generated videos and corresponding temporal attention maps. (2) We propose ByTheWay, which significantly improves the quality of T2V generation without necessitating training, introducing additional parameters, augmenting memory or sampling time. (3) ByTheWay can be seamlessly integrated with various mainstream open-source T2V backbones like AnimateDiff and VideoCrafter2, showing strong applicability and extensibility.
@@ 2.1 Text-to-Video Diffusion Models
Given a textual prompt, text-to-video (T2V) diffusion models [63, 64, 5, 6, 65, 17, 58] aim to synthesize image sequences that maintain both temporal consistency and textual alignment. Unlike text-to-image [66, 68, 67, 4] that emphasizes perfecting individual images, T2V poses a heightened challenge of maintaining both visual aesthetics for each frame and the realistic motion between frames. To this end, most approaches incorporate extra motion modeling modules into existing image diffusion architecture, leveraging the underlying image priors. For instance, AnimateDiff [9] introduced trainable temporal attention layers to frozen text-to-image models to effectively capture the frame-to-frame correlations. Some works [8, 7] combined temporal convolution modules and temporal attention layers for modeling short/long range dependencies. To alleviate motion synthesis difficulty, Ge et al. [59] suggested employing temporally related noise to enhance temporally consistent. Nevertheless, due to the scarcity of high-quality video data and the intricacies of motion synthesis, the current available T2V models still struggle to harmonize motion strength with motion consistency. This work identifies that the consistency across temporal attention blocks indicates the continuity of synthesized video sequences while the energy within the temporal attention maps dominates the magnitude of motion, and thus proposes a training-free strategy to unlock the potential of exiting T2V models by encouraging uniform motion modeling and enhanced frequency energy.
@@ 2.2 Diffusion Feature Control
Controlling diffusion features to manipulate specific attributes has been demonstrated to be an effective strategy in the realm of image and video synthesis [23, 61, 24, 26, 62]. Prompt2Prompt [22] revealed that the cross attention maps domain the image layout. DSG [53] proposed that spatial means of diffusion features represent the appearance, which offers simple approach for image property manipulation, such as size, shape, and location. FreeControl [27] suggested to perform image structure guidance by aligning the PCA features with given reference image in spatial self-attention block, providing a versatile counterpart of ControlNet [10]. DIFT [47] observed that the semantic correspondence can be directly extracted by spatially measuring the difference between diffusion feature. FreeU [52] suggested re-weighting the contribution of skip features and backbone features by using spectral modulation and structure-related scaling, promoting the emphasis on backbone semantics. In the field of video generation, MotionClone [51] demonstrated the sparse control of temporal attention maps facilitates a training-free motion transfer, enabling reference-based video generation. FreeInit[71] proposed to alleviate the initialization gap in video generation by iteratively refining the low-frequency components of initial latent, but suffers from increased inference cost and attenuated motion. UniCtrl[72] suggested to improve content alignment across frames by sharing the keys/values of first frame in self-attention layers, which produces reduced motion magnitude and thus requires extra branch for motion preservation. I4VGEN[73] decomposed text-to-video generation in a sequential manner, which demands the collaboration of I2V models and is characterized by higher inference cost. In this work, we propose Temporal Self-Guidance to facilitates uniform motion modeling across blocks by narrowing the disparities between temporal attention maps. This is work together with Fourier-based Motion Enhancement, which boosts motion magnitude by amplifying frequency energy, thus elevating the quality of the generated videos.
@@ 3.1 Latent Diffusion Model
In the context of T2V generation, latent diffusion model [56] is widely as backbone as its significant advancement in image synthesizing. Typically, based on a pre-trained autoencoder $\mathcal{E}(\cdot)$ and $\mathcal{D}(\cdot)$, video sequences are projected into the latent space, in which a denoising network $\epsilon_{\theta}$ is encouraged to learn the mapping from noised video latent $z_{t}$ to pure video latent $z_{0}$. Mathematically, the noised video latent $z_{t}$ obeys the following distribution:
@@
where $\bar{\alpha}_{t}$ is a pre-defined parameter representing noise schedule [29], $\epsilon\sim\mathcal{N}(0,1)$ is the added noise, and $t\sim\mathcal{U}(1,T)$ denotes time step. To restore $z_{0}$ from $z_{t}$, denoising network $\epsilon_{\theta}$ is forced to estimate the noise component in $z_{t}$, which can be expressed as:
@@
where $c$ represents the textual prompt. During sampling, $z_{t}$ is initialized with Gaussian noise and undergoes iterative denoising conditioned on $c$ for prompt-aligned generation.
Figure 2: Statistical patterns derived from T2V generation process. (a) Generated videos exhibiting structurally implausible and temporally inconsistent artifacts demonstrate greater disparity between the temporal attention maps of different decoder blocks. (b) After applying ByTheWay, the modeling disparity in original corrupted videos are reduced to the level of well-generated videos. (c) Videos with larger motion magnitude typically exhibit higher energy, in which the motion magnitude is measured by the estimated optical flow.
Figure 3: Temporal Self-Guidance. Temporal Self-Guidance contributes to the restoration of collapsed structures and consistency of motion in the generated video.
@@ 4.1 Temporal Self-Guidance
Temporal attention modules are extensively integrated at various hierarchical stages within the upsampling blocks of T2V architectures [8, 9, 6, 7]. These modules, derived from different tiers of the diffusion U-Net, are employed to capture inter-frame dependencies at multiple resolutions. We conjecture that, due to the limited capability in modeling motion, the nearby temporal attention maps struggle to capture large motion increments, which can lead to implausible structures and temporal inconsistencies. To substantiate this hypothesis, we analyzed 100 structurally and motion-degraded videos alongside 100 well-generated videos. The motion disparities are defined as the incremental differences in motion modeling across various blocks within the model, which are quantified by calculating the L2 difference between the temporal attention maps of the up_blocks.1 and the subsequent blocks. As illustrated in Fig. 2(a), it is observed that significant disparities between temporal attention maps across different blocks are associated with the occurrence of implausible structures and temporal inconsistencies.
To mitigate the excessive divergence between temporal attention maps across various upsampling blocks, we introduce a straightforward yet potent temporal self-guidance mechanism. This mechanism involves the infusion of the temporal attention map of up_blocks.1 into subsequent blocks, modulated by a guidance ratio $\alpha$. The adjustment can be mathematically modeled as:
@@
where $\mathcal{A}_{m}$ denotes the temporal attention map of $m$-th upsampling block ($m=2,3$), and $\mathcal{A}_{1}^{m}$ refers to the temporal attention map of up_blocks.1, which is upsampled to match the spatial dimensions of $\mathcal{A}_{m}$. As depicted in Fig. 2 (b) and Fig. 3, the implementation of temporal self-guidance effectively alleviates the excessive modeling disparity between different hierarchical levels of temporal attention modules, thereby diminishing the structurally implausible and temporally inconsistent artifacts in the resultant video generation.
@@ 4.2.1 Energy Representation of Motion Magnitude
Figure 4: Energy representation of video motion magnitude. Samples with richer motion typically exhibit a higher energy.
@@
$E=\frac{1}{F}\sum_{i=0}^{F-1}\sum_{j=0}^{F-1}||\mathcal{A}_{...,i,j}||^{2},\vspace{-0.5em}$ (4)
as illustrated in Fig. 4. To substantiate the correlation between the energy of the temporal attention map and the motion magnitude in the generated video, we employ the RAFT [55] to extract the optical flow, using the average magnitude of this flow as a metric for motion strength. Our findings reveal a positive correlation: videos with greater motion magnitudes are associated with higher spatially averaged energies within their temporal attention maps, as depicted in Fig. 2 (c). This insight motivates us to manipulate the motion magnitude in the generated videos by modulating the energy intensity of the temporal attention maps. By doing so, we aim to enhance the dynamism and variability of the motion in the videos.
4.2.2 Motion Enhancement by Frequency Re-weighting